# Disentangling the galactic and intergalactic components in 313 observed Lyman-alpha line profiles between redshift 0 and 5

### v1

@@ 4.8 Stacked Lyman-alpha line profile fitting
In Section 2 and 4, we have discussed the IGM+z and IGM-z models. These two models have very different training sets. On the one hand, IGM+z makes use of the IGM transmission curves from Byrohl & Gronke (2020) without the recalibration [^5] We refer here to the recalibration to the IGM transmission curves from Byrohl & Gronke (2020) so that they match the mean IGM redshift evolution by Faucher-Giguère et al. (2008). See ZP24. matching the redshift of the source.
@@
On the other hand, IGM-z uses the IGM transmission curves after recalibration to match the mean IGM transmission in Faucher-Giguère et al. (2008), and the IGM redshift is randomized. These differences cause the $f_{\rm esc}^{4\AA}$ distribution to evolve differently with redshift in the training sets (see section 3 in ZP24). Thus, the $f_{\rm esc}^{4\AA}$ predictions are different between these models. For example, the IGM-z model gives an unbiased $f_{\rm esc}^{4\AA}$ measurement from $z=0$ to $z=5.0$ for mocks with different $\langle f_{\rm esc}^{4\AA}\rangle$ redshift evolutions with $S/N_{p}$ =15.0 and $W_{\rm in}$ =0.1Å. Meanwhile, for the same spectral quality, IGM-z overpredicts $\langle f_{\rm esc}^{4\AA}\rangle$ by a 5% when $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.8$ at $z<1.0$, while at $z>1.0$ $\langle f_{\rm esc}^{4\AA}\rangle$ is recovered correctly (see section 4.3 in ZP24).
@@
Despite the differences between IGM+z and IGM-z, both models predict that the ${\rm Ly}\alpha$ line profiles at $z<0.5$ are attenuated by the IGM. IGM+z predicts a mean attenuation of $\sim 0.92$. Also, IGM-z predicts $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.89$. Both models predict individual values spawning from $\sim$0.8 to 1.0. If these lines are IGM attenuated, given the high signal-to-noise and resolution of HST ${\rm Ly}\alpha$ line profiles, a few individual absorbed systems should be visible.
@@
In Fig. 13, we show six examples of ${\rm Ly}\alpha$ line profiles that could be showing IGM features at $z<0.5$. Example A is a double peak ${\rm Ly}\alpha$ line profile that has a very good signal to noise in individual pixels. At $\Delta\lambda_{0}\sim-3\AA{}$, in the tail of the blue peak, there is an apparent absorption. In contrast, IGM+z, IGM-z, and NoIGM predict that the tail of the blue peak decreases smoothly. This kind of feature is not present in the shell model. Thus, this might be an IGM absorption. Moreover, in example A, the red tail is quite well fitted by the three models. However, the red peak observed spectrum falls faster than the models. This could indicate that the IGM is absorbing the core of the line. This is also visible in examples C and D.
@@
Also, in example B we find a plateau of flux $\sim 0$ at the core of the line (from $\Delta\lambda_{0}=\sim-1.5\AA{}$ to 0.3Å) [^6] Note that zELDA has problems reproducing these wide absorption droughts at line center since we restrict ourselves to $T=10^{4}\,$ gas. Higher (effective) temperatures lead to a wider absorption and thus can alleviate the need for central IGM absorption This is not the case for absorption features off line-center also marked in Fig. 13. . In comparison, the continuum of this source is detected around $0.05$. Additionally, this line shows a wide red peak that extends up to $\Delta\lambda_{0}\sim 4\AA{}$. Symmetrically, the blue peak seems to extend down to $\Delta\lambda_{0}\sim-4\AA{}$. As in case A, example B shows a decrease in flux in the blue peak at $\Delta\lambda_{0}\sim-3\AA{}$. Neither of these two features is found in the shell models and might be compatible with IGM absorption.
@@
Furthermore, in cases C, D and F, we find that the blue peak exhibits small drops in flux. However, considering the signal-to-noise of individual pixels, these drops are significant. The three models IGM+z, IGM-z, and NoIGM match well the red peak. The models and observations match the ascension of the blue peak ($\Delta\lambda_{0}$ between -0.5Å and 0.0Å) well – especially in cases C and D. Also, the predicted model is slightly above the observations, while they mat again at the tail of the blue peak ($\Delta\lambda_{0}$ between -2.0Å and -1.8Å).
@@
Finally, example E is again a double peak ${\rm Ly}\alpha$ line profile with very good signal-to-noise in individual pixels. We find that IGM+z and NoIGM models predict almost the same intrinsic line profile, while IGM+z predicts a line with the same red peak but with a more substantial blue peak. Focusing on the blue peak, we find that in the observed ${\rm Ly}\alpha$ line profile, there is more flux in the tail of the blue peak than in the IGM+z and NoIGM models. However, the blue tail is very well fitted by the IGM-z model. An absorbing system in the IGM could also cause this.
@@
In summary, across examples A–F, potential signatures of IGM absorption are consistently observed in the blue peak, including small flux drops and deviations from the smooth profiles predicted by the models. Additionally, features such as steeper declines in the red peaks and wider absorption plateaus at line center further suggest the impact of IGM on ${\rm Ly}\alpha$ line profiles. These findings highlight the IGM’s role in shaping ${\rm Ly}\alpha$ profiles, particularly in regions where the ‘shell model’ struggles to fully replicate the observed spectra [^7] Note that other models of intragalactic ${\rm Ly}\alpha$ radiative transfer will have similar problems to carve out the steep absorption features observed as radiative transfer and the associated frequency redistribution generally smooths spectra..
@@
Figure 14: Distribution of the difference between the input shell model properties and those recovered for the mock with HST-like quality. In the top panels $V_{\rm exp}$, $N_{\rm H}$ and $\tau_{a}$ are shown from left to right. In the bottom panels, $W_{\rm in}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA}$ are shown from left to right. The prediction using IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow, respectively. The dashed horizontal lines mark the distribution’s mean matching the used model’s color. The black dashed line is set at 0 for reference. In top right corner a table summing up the mean ($\mu$) and standard deviation ($\sigma$) is displayed for the three models.
@@
4.10 Possible biases when fitting an IGM attenuated line with a model without IGM
@@
In this section, we study the possible biases induced by fitting an IGM attenuated line with a model without IGM. For this goal, we use NoIGM as the model without IGM absorption. Meanwhile, we also compare the results for the explored models with IGM, IGM+z, and IGM-z.
@@
Regarding the MUSE-like sample (Fig. 15), the IGM+z and IGM-z models are generally less biased than NoIGM. In particular, IGM+z and IGM-z exhibit a bias smaller than 0.1 dex in $V_{\rm exp}$, $\tau_{a}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA}$. We find that IGM+z and IGM-z are biased 0.15 dex in $W_{\rm in}$. Moreover, we find that NoIGM is more biased than IGM+z and IGM-z, particularly in $V_{\rm exp}$ and $W_{\rm in}$. Interestingly, we find that NoIGM (0.15 dex) is slightly less biased than IGM+z and IGM-z ($\sim$0.23 dex) when estimating $N_{\rm H}$.
@@
5 Summary and conclusions
The latest version of zELDA includes different artificial neural network models to recover the galactic component from an IGM attenuated ${\rm Ly}\alpha$ line profile. These models are trained using mock ${\rm Ly}\alpha$ lines computed using grid of precomputed ’shell model’ (Gurung-López et al., 2022) with a full Monte Carlo radiative transfer treatment (LyaRT, Orsi et al., 2012) and the ${\rm Ly}\alpha$ IGM transmission published by Byrohl & Gronke (2020) based on IllustrisTNG100 simulation (Nelson et al., 2019).
  • In general, we find that ${\rm Ly}\alpha$ line profiles are little attenuated in HST data (z¡0.5). Meanwhile, in MUSE data ($z>3.0$), both, IGM+z and IGM-z, predict that the blue side of ${\rm Ly}\alpha$ is attenuated. The ISM ${\rm Ly}\alpha$ line profiles predicted by our models frequently exhibit a stronger blue peak than that observed. This trend becomes prominent at higher redshift. This suggests that ${\rm Ly}\alpha$ line profile reconstruction becomes more important at higher redshifts. Therefore, galaxy kinematic studies using the ${\rm Ly}\alpha$ line profile at high redshift might benefit from using the IGM reconstructed line profile rather than the observed line profile directly.
  • We quantify the IGM escape fraction, $f_{\rm esc}^{4\AA}$, for the HST and MUSE samples. We find that the mean $f_{\rm esc}^{4\AA}$ at $z<0.5$ is different depending on the ANN model. While IGM+z predicts $\langle f_{\rm esc}^{4\AA}\rangle$ =0.92, IGM-z predicts $\langle f_{\rm esc}^{4\AA}\rangle$ =0.89. As discussed, IGM+z tends to be biased towards 1.0 in this redshift range, so the measurement provided by IGM-z should be more reliable. Moreover, both ANN models agree well when determining $\langle f_{\rm esc}^{4\AA}\rangle$ in MUSE data. IGM+z and IGM-z predict that $\langle f_{\rm esc}^{4\AA}\rangle$ evolve from 0.87 at $z=3.0$ to 0.55 at $z=5.0$. This measurement can be helpful to obtain a measured ’IGM-free’ ${\rm Ly}\alpha$ luminosity function. Interestingly, our models predict that $z<0.5$ ${\rm Ly}\alpha$ line profiles can exhibit IGM features such as those of the ${\rm Ly}\alpha$ forest. We have shown six potential examples of this.
  • We have compared the mean IGM ${\rm Ly}\alpha$ escape fraction predicted by zELDA and the global ${\rm Ly}\alpha$ escape fraction. We find that $f_{\rm esc}^{4\AA}$ is higher than the global ${\rm Ly}\alpha$ escape fraction in the range where both quantities are measured. Extrapolating, we find that at $z=6$, the IGM escape fraction could match the global escape fraction. This would imply that the ISM escape fraction would be unity at $z>6$. Thus, at higher redshifts the global ${\rm Ly}\alpha$ escape fraction should match that of the IGM.
  • Next, we have analyzed the ${\rm Ly}\alpha$ stacked line profile. In the observed stacked line profile, a red peak that becomes slightly thinner with redshift is accompanied by a weaker blue peak that faints progressively at higher redshift. This is in good agreement with previous studies (Hayes et al., 2023). Interestingly, the IGM+z and IGM-z models predict that the stacked ${\rm Ly}\alpha$ line profile before entering the IGM does not evolve strongly with redshift. In the galactic spectrum, the blue peak amplitude is almost invariant, and the line profile becomes slightly thinner at higher redshift. This would imply that the mechanism that shapes the galactic ${\rm Ly}\alpha$ line profile is invariant with redshift, at least up to $z\sim 5$. Thus, the redshift evolution in the observed ${\rm Ly}\alpha$ line profiles can be attributed to the radiative transfer in the IGM.
  • We examined the biases that arise when fitting an IGM-attenuated ${\rm Ly}\alpha$ line profile using a model that does not account for the IGM component. Some shell model parameters, like the intrinsic equivalent width, $EW_{\rm in}$, can be accurately retrieved. However, parameters such as the neutral hydrogen column density, $N_{\rm H}$, and the bulk velocity, $V_{\rm exp}$, may be biased depending on the spectral quality. This highlights the importance of taking into account the IGM absorption in the ${\rm Ly}\alpha$ line profiles at high redshift.
Acknowledgements.
The authors acknowledge the financial support from the MICIU with funding from the European Union NextGenerationEU and Generalitat Valenciana in the call Programa de Planes Complementarios de I+D+i (PRTR 2022) Project (VAL-JPAS), reference ASFAE/2022/025. This work is part of the research Project PID2023-149420NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. This work is also supported by the project of excellence PROMETEO CIPROM/2023/21 of the Conselleria de Educación, Universidades y Empleo (Generalitat Valenciana). MG thanks the Max Planck Society for support through the Max Planck Research Group. DS acknowledges the support by the Tsinghua Shui Mu Scholarship, funding of the National Key R&D Program of China (grant no. 2023YFA1605600), the science research grants from the China Manned Space Project with no. CMS-CSST2021-A05, and the Tsinghua University Initiative Scientific Research Program (no. 20223080023). This research made use of matplotlib, a Python library for publication quality graphics (Hunter, 2007), NumPy (Harris et al., 2020) and SciPy (Virtanen et al., 2020).
Appendix A $f_{\rm esc}^{4\AA}$ inference by an alternative model

### v2

@@ 5.5 Comparison between the ${\rm Ly}\alpha$ global and IGM escape fractions
Interestingly, zELDA predicts that the ${\rm Ly}\alpha$ line profiles at $z<0.5$ are already attenuated by the IGM, while the mean IGM transmission is close to unity at $z=0$. In sections 2 and 4, we have discussed the IGM+z and IGM-z models. These two models have very different training sets. On the one hand, IGM+z makes use of the IGM transmission curves from Byrohl & Gronke (2020) without the recalibration [^7] We refer here to the recalibration to the IGM transmission curves from Byrohl & Gronke (2020) so that they match the mean IGM redshift evolution by Faucher-Giguère et al. (2008). See Z25. matching the redshift of the source. On the other hand, IGM-z uses the IGM transmission curves after recalibration to match the mean IGM transmission in Faucher-Giguère et al. (2008), and the IGM redshift is randomized. These differences cause the $f_{\rm esc}^{4\AA }$ distribution to evolve differently with redshift in the training sets (see section 3 in Z25). Thus, the $f_{\rm esc}^{4\AA }$ predictions are different between these models. For example, the IGM-z model gives an unbiased $f_{\rm esc}^{4\AA }$ measurement from $z=0$ to $z=5.0$ for mocks with different $\langle f_{\rm esc}^{4\AA }\rangle$ redshift evolutions with $S/N_{p}$ =15.0 and $W_{\rm in}$ =0.1Å. Meanwhile, for the same spectral quality, IGM-z overpredicts $\langle f_{\rm esc}^{4\AA }\rangle$ by a 5% when $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.8$ at $z<1.0$, while at $z>1.0$ $\langle f_{\rm esc}^{4\AA }\rangle$ is recovered correctly (see section 4.3 in Z25).
@@
Despite the differences between IGM+z and IGM-z, both models predict that the ${\rm Ly}\alpha$ line profiles at $z<0.5$ are attenuated by the IGM. IGM+z predicts a mean attenuation of $\sim 0.92$. Also, IGM-z predicts $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.89$. Both models predict individual values ranging from $\sim$0.8 to 1.0. If these lines are IGM attenuated, given the high signal-to-noise and resolution of HST ${\rm Ly}\alpha$ line profiles, a few individual absorbed systems should be visible, similar to the ${\rm Ly}\alpha$ forest.
@@
In Fig. 11, we show six examples of ${\rm Ly}\alpha$ line profiles that could be showing IGM features at $z<0.5$. Example A is a double peak ${\rm Ly}\alpha$ line profile that has a very good signal-to-noise in individual pixels. At $\Delta\lambda_{0}\sim-3\AA$ {}, in the tail of the blue peak, there is an apparent absorption. In contrast, IGM+z, IGM-z, and NoIGM predict that the tail of the blue peak decreases smoothly. This kind of feature is not present in the shell model. Thus, this might be an IGM absorption. Moreover, in example A, the red tail is quite well fitted by the three models. However, the red peak observed spectrum falls faster than the models. This could indicate that the IGM is absorbing the core of the line. This is also visible in examples C and D.
@@
Also, in example B we find a plateau of flux $\sim 0$ at the core of the line (from $\Delta\lambda_{0}=\sim-1.5\AA$ {} to 0.3Å) [^8] Note that zELDA has problems reproducing these wide absorption troughs at line center since we restrict ourselves to $T=10^{4}\,$K gas. Higher (effective) temperatures lead to a wider absorption and thus can alleviate the need for central IGM absorption. This is not the case for absorption features off-line-center also marked in Fig. 11. . In comparison, the continuum of this source is detected around $0.05$. Additionally, this line shows a wide red peak that extends up to $\Delta\lambda_{0}\sim 4\AA$ {}. Symmetrically, the blue peak seems to extend down to $\Delta\lambda_{0}\sim-4\AA$ {}. As in case A, example B shows a decrease in flux in the blue peak at $\Delta\lambda_{0}\sim-3\AA$ {}. Neither of these two features is found in the shell models and might be compatible with IGM absorption.
@@
Furthermore, in cases C, D and F, we find that the blue peak exhibits small drops in flux. However, considering the signal-to-noise of individual pixels, these drops are significant. The three models IGM+z, IGM-z, and NoIGM match well the red peak. The models and observations match the ascension of the blue peak ($\Delta\lambda_{0}$ between -0.5Å and 0.0Å) well – especially in cases C and D. Also, the predicted model is slightly above the observations, while they match again at the tail of the blue peak ($\Delta\lambda_{0}$ between -2.0Å and -1.8Å).
@@
Finally, example E is a double peak ${\rm Ly}\alpha$ line profile with very good signal-to-noise in individual pixels. We find that IGM+z and NoIGM models predict almost the same intrinsic line profile, while IGM+z predicts a line with the same red peak but with a more substantial blue peak. Focusing on the blue peak, we find that in the observed ${\rm Ly}\alpha$ line profile, there is more flux in the tail of the blue peak than in the IGM+z and NoIGM models. However, the blue tail is very well fitted by the IGM-z model. An absorbing system in the IGM could also cause this.
@@
In summary, across examples A–F, potential signatures of IGM absorption are consistently observed in the blue peak, including small flux drops and deviations from the smooth profiles predicted by our shell models. While more sophisticated intragalactic ${\rm Ly}\alpha$ radiative transfer models might provide somewhat better fits without IGM attenuation, note that generally such models have similar problems in carving out steep absorption features due to the typical diffusive frequency redistribution smoothing spectra. Additionally, features such as steeper declines in the red peaks and wider absorption plateaus at line center could further hint at the impact of IGM on ${\rm Ly}\alpha$ line profiles. Hence, these findings highlight the IGM’s potential role in shaping ${\rm Ly}\alpha$ profiles down to low redshifts, particularly in regions where intragalactic radiative transfer models struggle to fully replicate the observed spectra.
6 Summary and conclusions
The latest version of zELDA includes different artificial neural network models to recover the galactic component from an IGM attenuated ${\rm Ly}\alpha$ line profile. These models are trained using mock ${\rm Ly}\alpha$ lines computed using a grid of precomputed ’shell model’ (Gurung-López et al., 2022) with a full Monte Carlo radiative transfer treatment (LyaRT, Orsi et al., 2012) and the ${\rm Ly}\alpha$ IGM transmission published by Byrohl & Gronke (2020) based on TNG100 simulation (Nelson et al., 2019).
  • In general, we find that ${\rm Ly}\alpha$ line profiles are little attenuated in HST data (z<<0.5). Meanwhile, in MUSE data ($z>3.0$), both, IGM+z and IGM-z, predict that the blue side of ${\rm Ly}\alpha$ is attenuated. The ISM ${\rm Ly}\alpha$ line profiles predicted by our models frequently exhibit a stronger blue peak than that observed. This trend becomes prominent at higher redshifts. This suggests that ${\rm Ly}\alpha$ line profile reconstruction becomes more important at higher redshifts. Therefore, galaxy kinematic studies using the ${\rm Ly}\alpha$ line profile at high redshift might benefit from using the IGM reconstructed line profile rather than the observed line profile directly.
  • We quantify the ${\rm Ly}\alpha$ line profile IGM escape fraction, $f_{\rm esc}^{4\AA }$, for the HST and MUSE samples. We find that the mean $f_{\rm esc}^{4\AA }$ at $z<0.5$ is different depending on the ANN model. While IGM+z predicts $\langle f_{\rm esc}^{4\AA }\rangle$ =0.92, IGM-z predicts $\langle f_{\rm esc}^{4\AA }\rangle$ =0.89. As discussed, IGM+z tends to be biased towards 1.0 in this redshift range, so the measurement provided by IGM-z should be more reliable. Moreover, both ANN models agree well when determining $\langle f_{\rm esc}^{4\AA }\rangle$ in MUSE data. IGM+z and IGM-z predict that $\langle f_{\rm esc}^{4\AA }\rangle$ evolve from 0.87 at $z=3.0$ to 0.55 at $z=5.0$. This measurement can be helpful to obtain a measured ’IGM-free’ ${\rm Ly}\alpha$ luminosity function. Interestingly, our models predict that $z<0.5$ ${\rm Ly}\alpha$ line profiles can exhibit IGM features such as those of the ${\rm Ly}\alpha$ forest. We have shown six potential examples of this.
  • The validation analyses confirm that zELDA provides robust reconstructions of ${\rm Ly}\alpha$ line profiles and reliable estimates of the redshift evolution of the IGM escape fraction. The method achieves sub-angstrom redshift accuracy ($\sim$0.3Å). The reconstructed profiles reproduce the red tail of the emission line with high fidelity and, when stacked, show little intrinsic evolution with redshift, indicating that the observed spectral variations are primarily driven by IGM absorption, in agreement with previous works (Hayes et al., 2021a).
  • We find that our estimate of the ${\rm Ly}\alpha$ IGM escape fraction is consistent with both the mean IGM optical depth and the global ${\rm Ly}\alpha$ escape fraction reported in the literature. Our results suggest that at $z\sim 5$ the IGM is the main absorber of the ${\rm Ly}\alpha$ emission line, attenuating $\sim 60\%$ of the total flux, while the combined contribution of the CGM and ISM accounts for the remaining $\sim 40\%$. If the observed trend of a decreasing $\langle f_{\rm esc}^{4\AA }\rangle$ with redshift extends to higher redshifts, the IGM would become even more dominant than the ISM/CGM in determining the observability of ${\rm Ly}\alpha$ emitters.
Acknowledgements
The authors acknowledge the financial support from the MICIU with funding from the European Union NextGenerationEU and Generalitat Valenciana in the call Programa de Planes Complementarios de I+D+i (PRTR 2022) Project (VAL-JPAS), reference ASFAE/2022/025. This work is part of the research Project PID2023-149420NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. This work is also supported by the project of excellence PROMETEO CIPROM/2023/21 of the Conselleria de Educación, Universidades y Empleo (Generalitat Valenciana). MG thanks the Max Planck Society for support through the Max Planck Research Group, and the European Union for support through ERC-2024-STG 101165038 (ReMMU). DS acknowledges the support by the Tsinghua Shui Mu Scholarship, funding of the National Key R&D Program of China (grant no. 2023YFA1605600), the science research grants from the China Manned Space Project with no. CMS-CSST2021-A05, and the Tsinghua University Initiative Scientific Research Program (no. 20223080023).
Appendix A Distribution of outflow parameters for IGM+z and IGM-z
Figure 12: Comparison of the ANN output and LASD variables in the observed ${\rm Ly}\alpha$ line profiles predicted by the same models. $EW_{\rm obs}$ and $L_{\rm Ly\alpha}$ were obtained from LASD. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$, $EW_{\rm in}$, $EW_{\rm ob}$ and $W_{\rm in}$ in Å and $L_{\rm Ly\alpha}$ in $erg/s$. The blue dots show the values predicted by IGM+z while the green dots show those of IGM-z. The grey dots show the properties that include measurements by LASD. The histograms on the diagonal show the distribution of the parameters in the corresponding color for each model. The strength of the correlations is quantified in Fig. 3
Fig. 12 shows the predicted parameters plotted against each other: $V_{\rm exp}$, $N_{\rm H}$, $\tau_{a}$, $EW_{\rm in}$, $W_{\rm in}$, $z$, $f_{\rm esc}^{4\AA }$, $EW_{\rm ob}$, and $L_{\rm Ly\alpha}$, arranged from left to right and top to bottom. The upper-right triangle (green and grey) corresponds to the properties predicted by IGM-z, while the lower-left triangle (blue and grey) corresponds to IGM+z. The diagonal panels show the one-dimensional distributions of each parameter. The strength of the (anti)correlations is quantified in Fig. 3 and discussed in Sect.4.2.
Appendix B Comparison of Outflow Properties across Models
In Fig. 14, we show a comparison between the output variables of IGM+z (X-axis) and those predicted by NoIGM (Y-axis). In general, we find that there is a Spearman correlation coefficient higher than $\sim 0.7$ for the whole sample. The Spearman correlation coefficient is greater in the low redshift sample than in the high redshift sample. For example, for $V_{\rm exp}$, $\rho_{z<3}=0.72$ while $\rho_{z>3}=0.69$ and for $N_{\rm H}$ $\rho_{z<3}=0.83$ while $\rho_{z>3}=0.64$. The IGM+z model predicts higher values for $V_{\rm exp}$ at high redshift. Meanwhile, IGM+z predicts lower values of $W_{\rm in}$ than NoIGM at high redshift.
Appendix C Possible biases when fitting an IGM attenuated line with a model without IGM
@@
Figure 15: Distribution of the difference between the input shell model properties and those recovered for the mock with HST-like quality. In the top panels $V_{\rm exp}$, $N_{\rm H}$ and $\tau_{a}$ are shown from left to right. In the bottom panels, $W_{\rm in}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA }$ are shown from left to right. The prediction using IGM+z, IGM-z, and NoIGM is shown in blue, green, and yellow, respectively. The dashed horizontal lines indicate the distribution’s mean, matching the color used in the model. The black dashed line is set at 0 for reference. In top right corner a table summing up the mean ($\mu$) and standard deviation ($\sigma$) is displayed for the three models.
@@
In this section, we study the possible biases induced by fitting an IGM-attenuated line with a model without IGM. For this goal, we use NoIGM as the model without IGM absorption. Meanwhile, we also compare the results for the explored models with IGM, IGM+z, and IGM-z.
@@
Regarding the MUSE-like sample (Fig. 16), the IGM+z and IGM-z models are generally less biased than NoIGM. In particular, IGM+z and IGM-z exhibit a bias smaller than 0.1 dex in $V_{\rm exp}$, $\tau_{a}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA }$. We find that IGM+z and IGM-z are biased by $0.15$ dex in $W_{\rm in}$. Moreover, we find that NoIGM is more biased than IGM+z and IGM-z, particularly in $V_{\rm exp}$ and $W_{\rm in}$. Interestingly, we find that NoIGM (0.15 dex) is slightly less biased than IGM+z and IGM-z ($\sim$0.23 dex) when estimating $N_{\rm H}$.
@@
Appendix D Property correlation in MUSE and HST
In Fig.17 we show the Spearman correlation coefficients for the HST (left) and MUSE (right) samples. We find that in general, all the correlations found in the complete sample (see Fig.3) are also present in the individual samples, such as the correlation between $EW_{\rm in}$ and $EW_{\rm ob}$.
Other correlations existing in the combined sample are only present in one of them. For example, the anti-correlation between $f_{\rm esc}^{4\AA }$ and redshift is present in the MUSE and combined sample. However, no trend is found within the HST sources.
Appendix E $f_{\rm esc}^{4\AA }$ inference by an alternative model